# Thermodynamic Basis of Sugar-Dependent Polymer Stabilization: Informing Biologic Formulation Design

### v1

@@ 3.4 Sugar-Specific Stabilization Mechanisms: The Role of Interactions and Entropy
The unfolding free energies, shown in Fig. 2e, indicate that $\beta$-fructose stabilizes the HP more effectively than $\alpha$-glucose at equivalent concentrations. To understand the origin of this difference, we compare the thermodynamic decomposition of the PMF obtained for the two sugars. The results of this decomposition for all concentrations are shown in SI Fig. S5. We group the polymer-sugar interactions and polymer-water interactions into a single polymer-solvent interaction term ($\Delta\Delta E_{p-solv}$) as per Eq. 6. At low concentrations (up to 1 M), $\Delta\Delta E_{p-solv}$ more strongly favor folding in $\beta$-fructose than in $\alpha$-glucose solution (SI Fig. S5). However, at 2 M, $\Delta\Delta E_{p-solv}$ favored folding more in $\alpha$-glucose than in $\beta$-fructose, although the difference was not statistically significant. Despite this, $\beta$-fructose exhibited a higher overall $\Delta\Delta G_{u}$, indicating greater stabilization of the folded state (Fig. 7). This difference is reconciled by considering the $\Delta\Delta G_{rem}$ term, which favored folding more in $\beta$-fructose than in $\alpha$-glucose. Together, these findings illustrate a concentration-dependent stabilization mechanism, wherein polymer-solvent interaction energies differentiate stabilizing excipients at low concentrations, while entropic effects become increasingly important at higher concentrations.
@@ 3.5 Local Mixing Entropy as a Proxy for Entropic Contributions
This metric is evaluated by defining a local solvent shell around the polymer, consistent with the region used to compute preferential interaction coefficients. Within this shell, we calculate the atom number fractions of water and sugar molecules in each simulation frame. Based on these fractions, we compute the local mixing entropy as described in Section 2.5, and determine the entropy difference between the folded and unfolded states, denoted as $\Delta\Delta S_{mix}=\Delta S_{mix,unfolded}-\Delta S_{mix,folded}$. Positive values of $\Delta\Delta S_{mix}$ indicate greater solvent compositional heterogeneity—and thus higher configurational entropy—in the unfolded state.
@@
These observations build upon previous studies, which have primarily attributed protein stability to excluded volume effects and specific protein-sugar interactions 43, 42, 54. Our findings highlight polymer-solvent interaction entropy as an additional, previously underexplored strategy for enhancing stability in biological formulations. The interplay of direct interactions, excluded volume, and local configurational entropy collectively governs the stabilizing capacity of sugar solutions, with different contributions being decisive at different concentrations. Future studies that systematically manipulate polymer-solvent interaction entropy - through sugar mixtures, excipient design, or solvent composition - will be key to determining its broader applicability in optimizing sugar-based excipients for biological formulations.
@@ Conclusions
Overall, our results underscore the complexity of protein stabilization mechanisms in sugar solutions and emphasize the importance of considering concentration-dependent effects, the nature of sugar-protein interactions, and the protein-solvent entropy when designing formulations. These findings contribute to a more nuanced understanding of excipient-protein interactions, which can inform more effective formulation strategies in biopharmaceutical design. Future work will aim to extend these findings to protein systems with diverse surface chemistries and folding behaviors to develop a more generalized understanding of sugar-mediated stabilization mechanisms.
@@
This material is based upon work supported by the National Science Foundation under DMREF Grant No. 2118788, 2118693, and 2118638. The Minnesota Supercomputing Institute at the University of Minnesota Twin Cities provided computational resources.

### v2

@@ 3.4 Sugar-Specific Stabilization Mechanisms: The Role of Interactions and Entropy
The unfolding free energies, shown in Fig. 2e, indicate that $\beta$-fructose promotes folding of HP more effectively than $\alpha$-glucose at equivalent concentrations (0.99 vs 0.74 kT at 0.25 M, 1.41 vs 1.24 kT at 0.5 M, 2.66 vs 1.81 kT at 1 M, and 0.66 vs 0.44 kT at 2 M). To understand the origin of this difference, we compare the thermodynamic decomposition of the PMF obtained for the two sugars. The results of this decomposition for all concentrations are shown in SI Fig. S5. We group the polymer-sugar interactions and polymer-water interactions into a single polymer-solvent interaction term ($\Delta\Delta E_{p-solv}$) as per Eq. 6. At low concentrations (up to 1 M), $\Delta\Delta E_{p-solv}$ more strongly favor folding in $\beta$-fructose than in $\alpha$-glucose solution (SI Fig. S5). However, at 2 M, $\Delta\Delta E_{p-solv}$ favored folding more in $\alpha$-glucose than in $\beta$-fructose, although the difference was not statistically significant. Despite this, $\beta$-fructose exhibited a higher overall $\Delta\Delta G_{u}$, indicating greater stabilization of the folded state (Fig. 7). This difference is reconciled by considering the $\Delta\Delta G_{rem}$ term, which favored folding more in $\beta$-fructose than in $\alpha$-glucose. Together, these findings illustrate a concentration-dependent stabilization mechanism, wherein polymer-solvent interaction energies differentiate stabilizing excipients at low concentrations, while entropic effects become increasingly important at higher concentrations.
@@ 3.5 Local Mixing Entropy as a Proxy for Entropic Contributions
This metric is evaluated by defining a local solvent shell around the polymer, consistent with the region used to compute preferential interaction coefficients. Within this shell, we calculate the atom number fractions of water and sugar atoms in each simulation frame, ensuring that each atom is uniquely identified and counted only once, even in geometrically overlapping shell regions. Because these fractions are normalized, $\Delta\Delta S_{mix}$ is independent of the total shell volume or the number of local atoms, reflecting only changes in the local compositional diversity upon folding. Based on these fractions, we compute the local mixing entropy as described in Section 2.5, and determine the entropy difference between the folded and unfolded states, denoted as $\Delta\Delta S_{mix}=\Delta S_{mix,unfolded}-\Delta S_{mix,folded}$. Positive values of $\Delta\Delta S_{mix}$ indicate greater solvent compositional heterogeneity—and thus higher configurational entropy—in the unfolded state.
In our thermodynamic decomposition, $\Delta S_{ex}$ represents the entropy change associated with depletion due to excluded volume, and $\Delta S_{si}$ corresponds to configuration-specific interaction effects. In contrast, $\Delta\Delta S_{mix}$ captures compositional changes in the local solvent environment, which may arise from either excluded-volume–driven depletion or specific sugar–polymer interactions. Therefore, this metric quantifies the compositional entropy component of polymer stabilization, rather providing a measure of the number of local atoms or the total configurational entropy.
@@
We further analyzed $S_{mix}$ as a function of the polymer radius of gyration ($R_{g}$) to examine the local compositional entropy around the stable polymer conformations. This analysis revealed that the hairpin configuration (state II) exhibits $S_{mix}$ values comparable to or slightly higher than those of the fully collapsed state (state I). At low concentrations (sugar atom number fraction $x_{s}$ < 0.5), this indicates similar or greater local sugar enrichment in the hairpin state. At higher concentrations ($x_{s}$ > 0.5), the $x_{s}$ (state I) > $x_{s}$ (state II) resulting in higher $S_{mix}$ in State II. This happens because at higher concentrations, with the increase in $R_{g}$, while the number of sugar atoms in the vicinity of the polymer increases, the increase in water atoms is higher. We hypothesize that this could be related to the smaller size of water molecules being able to penetrate the already crowded solvent shell resulting in an increase in mixing entropy. Representative $S_{mix}$$R_{g}$, $x_{s}$-$R_{g}$, $n_{s}$-$R_{g}$, and $n_{w}$-$R_{g}$ plots illustrating these trends are provided in the SI Fig. S10.
Overall, $\Delta\Delta S_{mix}$ captures essential features of polymer-solvent interaction entropy and offers a physically interpretable and computationally efficient measure to assess the entropic component of polymer stabilization. These observations build upon previous studies, which have primarily attributed protein stability to excluded volume effects and specific protein-sugar interactions 43, 42, 54. Our findings highlight polymer-solvent interaction entropy as an additional, previously underexplored strategy for enhancing stability in biological formulations. The interplay of direct interactions, excluded volume, and local configurational entropy collectively governs the stabilizing capacity of sugar solutions, with different contributions being decisive at different concentrations. Future studies that systematically manipulate polymer-solvent interaction entropy - through sugar mixtures, excipient design, or solvent composition - will be key to determining its broader applicability in optimizing sugar-based excipients for biological formulations.
@@ Conclusions
Overall, our results underscore the complexity of protein stabilization mechanisms in sugar solutions and emphasize the importance of considering concentration-dependent effects, the nature of sugar-protein interactions, and the protein-solvent entropy when designing formulations. These findings contribute to a more nuanced understanding of excipient-protein interactions, which can inform more effective formulation strategies in biopharmaceutical design. Additionally, the framework presented here is also relevant to understanding crowding effects on protein folding equilibria in cellular environments. While the present work focused on the (++––) charge sequence, exploring alternative arrangements could further elucidate the influence of charge distribution on polymer folding and excipient-mediated stabilization. Future work will also aim to extend these findings to protein systems with diverse surface chemistries and folding behaviors to develop a more generalized understanding of sugar-mediated stabilization mechanisms.
4 Supporting Information
Supporting information includes KBP force field validation results, additional simulation details, convergence check for PMFs, variation of preferential interaction coefficient with local domain cutoff radius, unfolding free energy vs change in preferential interactions, PMF decomposition and subgroup interaction energies, radial distribution function of $\beta$-fructose $\mathrm{CH_{2}OH}$ groups around the HP, change in mixing entropy with local radius, PMF decomposition results with replicate runs.
@@ 4 Supporting Information
This material is based upon work supported by the National Science Foundation under DMREF Grant No. 2325392, 2118788, 2118693, and 2118638. The Minnesota Supercomputing Institute at the University of Minnesota Twin Cities provided computational resources.